OpenStudy (anonymous):

Adele is 5 years older than Timothy. In three years, Timothy will be 2/3 of Adele’s age. What is Adele’s current age?

OpenStudy (anonymous):

A=T+5, T=2/3 A Plug the second value of T into the first equation, then subtract 3.

OpenStudy (anonymous):

She's 12?

OpenStudy (anonymous):

aish I'm confusing myself >.<

OpenStudy (phi):

ok on the first equation A= T+5 for the second equation In three years, Timothy will be 2/3 of Adele’s age. in three years, Timothy will be (T+3) years old and Adelle will be (A+3) so you should write (T+3)= 2/3 * (A+3)

OpenStudy (anonymous):

Yup. Good job!